OpenStudy (anonymous):

A cup of rolled oats (x) provided 310 calories. A cup of wheat flakes (y) provided 290 calories. A new breakfast cereal combines wheat and oats to provide 302 calories per cup. How much of each grain does 1 cup of the cereal contain? a.. Write one equation about the amount of grain b.Write a second equation about the number of calories.

OpenStudy (anonymous):

a) We use x cups of oats and (1-x) cups of flakes to get one cup of the new combined cereal. 310*x + 290 * (1-x) = 302 Solving this you'll get: x = 0.6 = 3/5 1-x = 0.4 = 2/5 3/5 of oats and 2/5 of flakes will give you the perfect cereal :)

OpenStudy (anonymous):

what are the equations?

OpenStudy (anonymous):

310*x + 290 * y = 302 x+y=1

OpenStudy (anonymous):

:) Marina thank u